OpenStudy (anonymous):

how to solve this e^(x+1) = 26

jimthompson5910 (jim_thompson5910):

First convert to logarithmic form. What do you get when you do this

OpenStudy (anonymous):

log^26=x+1 ?

jimthompson5910 (jim_thompson5910):

close, it's more like log(e, 26) = x+1 or ln(26) = x+1

OpenStudy (anonymous):

okay so ln(26) = 3.26 right

OpenStudy (anonymous):

so then i just move 1 to the other side

OpenStudy (anonymous):

is it x=2.26

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (precal):